OpenStudy (anonymous):

If the graph of the given equation is a circle, find its center and radius. If the equation has no graph, say so.

OpenStudy (anonymous):

\[x^2+y^2=-8y\]

OpenStudy (anonymous):

@terenzreignz

terenzreignz (terenzreignz):

Ahh... the joys of analytic geometry. More like woes, in your case, but that's why I'm here, to turn them into joys ^_^ Okay, every equation of a circle can be written as follows, yes? \[\Large (x-\color{red}h)^2 + (y-\color{green}k)^2 = \color{blue}r^2\] Where \(\large (\color{red}h,\color{green}k) \) is the centre, and \(\large \color{blue} r\) is the radius.

OpenStudy (anonymous):

yes so I had so far.... \[x^2 + (y^2+8y+?)=0\]

terenzreignz (terenzreignz):

Very good, <what was your name again?>. Now, I need you to possess a certain skill... it's called "completing the square" Have you heard of it? ^_^

OpenStudy (anonymous):

hello! hahaha, yes but I kind of have trouble with it, or need to refresh my memory...

terenzreignz (terenzreignz):

Completing the Square 101 Imagine you have something of the form \[\Large z^2 + \color{purple}pz = k\] And for some reason or another, you need to express it in terms of a perfect square trinomial. The first thing to watch out for is to make sure that the square bit... (this part)\[\Large \color{red}{z^2 }+ \color{purple}pz = k\] is unadorned... meaning, it has no number next to it... if it does, well, remedy it by dividing the entire expression by that number... So now, pay attention to the coefficient of the 'unsquared' variable, in this case, \(\large \color{purple}p\). What you need to do is halve it, resulting in \(\Large \frac{\color{purple}p}{2}\) and then square it, yielding \(\Large \frac{\color{purple}p^2}{4}\). Next, you add this to both sides of your equation, like so.. \[\Large z^2 + \color{purple}pz+\frac{\color{purple}p^2}4 = k+\frac{\color{purple}p^2}{4}\] You'll notice that the left side is now a perfect square, in particular, the square of \(\Large z+\frac{\color{purple}p}2\)... \[\Large \left(z+\frac{\color{purple}p}2\right)^2=k+\frac{\color{purple}p^2}{4}\] And that's it! You have successfully 'completed the square' :)

terenzreignz (terenzreignz):

Read and understand, and then I want you to complete the square of the part \[\Large y^2 + 8y\]

terenzreignz (terenzreignz):

And if you DON'T understand, tell me ASAP :)

OpenStudy (anonymous):

okay hold on for a bit.

OpenStudy (anonymous):

so why do I halve P before I square it, why not just square it?

terenzreignz (terenzreignz):

Good question. It's because a perfect square trinomial, say, we square a + qb it'd look like \[\Large (a+qb)^2 = a^2 + \color{orange}{2q}ab + \color{green}{q^2}b^2\] And notice that the orange coefficient is just twice the square root of green coefficient or equivalently The green coefficient is the square of HALF the orange coefficient. Clear? ^_^

OpenStudy (anonymous):

okay haha it's a little bit confusing but I think I get it.

terenzreignz (terenzreignz):

Now, get back to work >:)

terenzreignz (terenzreignz):

Complete the square in \[\Large y^2 + 8y\] in case you get lost :P

OpenStudy (anonymous):

okay one sec

OpenStudy (anonymous):

so then i get \[y^2 + 8y+16\]

terenzreignz (terenzreignz):

Great ^_^ So do we just do this? \[\Large x^2 + y^2 + 8y \color{blue}{+16}=0\]?

OpenStudy (anonymous):

yes.. and we group so, \[x^2 + (y^2+8y+16)=0 \]

terenzreignz (terenzreignz):

Nope. :P

terenzreignz (terenzreignz):

Stop right there and LISTEN. (first) ^_^

terenzreignz (terenzreignz):

We can't just do this. \[\Large x^2 + y^2 + 8y \color{red}{+ 16 }= 0\] It's not 'fair' to the equation. We just added 16 to the left-side, we have to do it to the right-side too :P \[\Large x^2 + y^2 + 8y \color{red}{+16}= 0 \color{blue}{+16}\] Understood? ^_^

OpenStudy (anonymous):

yes hahaha That was what I was gonna do next! :D

OpenStudy (anonymous):

I was supposed to do that first then group... lol oops

terenzreignz (terenzreignz):

Could have... would have... didn't :P Anyway, proceed as thus.

OpenStudy (anonymous):

\[x^2+(y+4)^2=16\]

terenzreignz (terenzreignz):

Impressive. You learn quickly ^_^ Now compare with \[\Large (x-\color{red}h)^2 + (y-\color{green}k)^2 = \color{blue}r^2\] And determine your values for \(\color{red}h\), \(\color{green}k\), and \(\color{blue}r\).

OpenStudy (anonymous):

which means C=(0,-4) and R= 4 ?

OpenStudy (anonymous):

h is 0, k is -4 r=4

terenzreignz (terenzreignz):

What else could I say? Well done ^_^

OpenStudy (anonymous):

Awesome! Thanks again! might need more help in the near future with new material lol... if you're available... I'll tag you! :D

terenzreignz (terenzreignz):

Bloody spoiled ..... >:( ...lol JK. Better hope I'm still online by then. I *am* getting sleepy :P